OpenStudy (dricohernandez):

Two identical closely spaced circular disks form a parallel-plate capacitor. Transferring 2.2×109 electrons from one disk to the other causes the electric field strength between them to be 1.3×105 N/C .What are the diameters of the disks?

OpenStudy (irishboy123):

what formulae do you know?? i can roll with \(C = \dfrac{\epsilon_o A}{d} = \dfrac{Q}{V}\) \( A = \dfrac{Q d}{\epsilon_o V}\) And the electric field in parallel plates is \(E = \dfrac{ V}{ d}\) So.... \( A = \dfrac{Q }{\epsilon_o E}\) you've got those numbers, i reckon

OpenStudy (dricohernandez):

Q/4πεod²

OpenStudy (irishboy123):

\(D = \sqrt{ \dfrac{4Q}{\pi \epsilon_o E} } \)

OpenStudy (irishboy123):

As in \(A = \dfrac{\pi D^2}{4}\)

OpenStudy (dricohernandez):

so what would the answer be?

OpenStudy (irishboy123):

no idea mate, i'd need to find a calculator